OpenStudy (anonymous):

Solve by substitute { 5x+4y=12 {3x+4y=4 Please help

OpenStudy (texaschic101):

5x + 4y = 12 4y = -5x + 12 y = -5/4x + 3 now sub -5/4x + 3 in for y into the other equation 3x + 4y = 4 3x + 4(-5/4x + 3) = 4 3x - 5x + 12 = 4 3x - 5x = 4 - 12 -2x = - 8 x = 4 now sub 4 in for x in either equation to find y 5x + 4y = 12 5(4) + 4y = 12 20 + 4y = 12 4y = 12 - 20 4y = - 8 y = -2 now we check.. 3x + 4y = 4 3(4) + 4(-2) = 4 12 - 8 = 4 4 = 4 (correct) answer is : x = 4 and y = -2 or (4,-2)